can you solve these problems out in step?
solve and graph on a number line.
5-3x<14
solve x
x+2 = 3/7

a brownie recipe that makes 36 brownies calls for 1 1/2 cups of sugar. How many cups of sugar are needed to make 24 brownies?

Answer by xcentaur(357)

solve and graph on a number line.
5-3x<14


-3x<14-5
-3x<9
dividing by negative number reverses inequality
x>-3


solve x
x+2 = 3/7

x=(3/7)-2
x=(3-14)/7=(-11/7)


a brownie recipe that makes 36 brownies calls for 1 1/2 cups of sugar. How many cups of sugar are needed to make 24 brownies?
36 - 1.5
24 - ?
x=(24*1.5)/36
=(8*1.5)/12
=12/12
=1 cup


Prabhat.